# Resisting Backdoor Attacks in Federated Learning via Bidirectional Elections and Individual Perspective

### v1

@@ Abstract
Existing approaches defend against backdoor attacks in federated learning (FL) mainly through a) mitigating the impact of infected models, or b) excluding infected models. The former negatively impacts model accuracy, while the latter usually relies on globally clear boundaries between benign and infected model updates. However, model updates are easy to be mixed and scattered throughout in reality due to the diverse distributions of local data. This work focuses on excluding infected models in FL. Unlike previous perspectives from a global view, we propose Snowball, a novel anti-backdoor FL framework through bidirectional elections from an individual perspective inspired by one principle deduced by us and two principles in FL and deep learning. It is characterized by a) bottom-up election, where each candidate model update votes to several peer ones such that a few model updates are elected as selectees for aggregation; and b) top-down election, where selectees progressively enlarge themselves through picking up from the candidates. We compare Snowball with state-of-the-art defenses to backdoor attacks in FL on five real-world datasets, demonstrating its superior resistance to backdoor attacks and slight impact on the accuracy of the global model.
@@ 1 Introduction
Existing defenses to backdoor attacks in FL are mainly based on a) mitigating the impact of infected models (Bagdasaryan et al. 2020; Sun et al. 2019; Xie et al. 2021; Nguyen et al. 2022; Zhang et al. 2023) or b) excluding infected models based on their deviations (Blanchard et al. 2017; Ozdayi, Kantarcioglu, and Gel 2021; Fung, Yoon, and Beschastnikh 2018; Rieger et al. 2022; Li et al. 2020a; Shejwalkar and Houmansadr 2021; Zhang et al. 2022; Shi et al. 2022; Nguyen et al. 2022). The former may negatively impact global model accuracy (Yu et al. 2021). The latter assumes globally clear boundaries between benign and infected model updates (Zeng et al. 2022). However, backdoor attacks typically manipulate a limited subset of parameters, resulting in similarity among benign and infected model updates. Besides, the nature of Non-Independent and Identically Distributed (non-IID) data in FL increases diversity among model updates.
@@
Actually, benign and infected model updates are easy to be mixed with complicatedly non-IID data (practical non-IID (Hsu, Qi, and Brown 2019; Huang et al. 2021) and feature distribution skew (Tan et al. 2022)), or with not very high poison data ratio (PDR). We experimentally demonstrate it in Figure 1, where benign and infected updates are mixedly scattered. In such cases, anomaly detection techniques based on linear similarity may not perform satisfactorily. Besides, when facing relatively high malicious client ratio (MCR), infected model updates are easier to be mistreated as benign ones, however, many existing defenses are only evaluated with MCR $\leq$ 10% (Xie et al. 2021; Ozdayi, Kantarcioglu, and Gel 2021; Zeng et al. 2022; Lu et al. 2022). Although model deviations may be better captured through nonlinear neural networks, the patterns of benign models in FL are usually hard to be acquired due to unpredictable distributions and trajectory shifts of model updates. (Li et al. 2020a) use the test data to generate model weights for training the detection model, but the test data with similar distribution to all clients may usually be unavailable on the server. Besides, model weights follow extremely complex distributions, making them very difficult to learn.
To better leverage powerful neural networks to detect malicious models, we propose Snowball, an anti-backdoor FL framework taking advantages of linear and non-linear approaches, i.e, without need of pre-defined benign patterns and the powerful capability to capture model deviations, respectively. It treats each model update an agent electing model updates for aggregation with an individual perspective, where the motivation comes from that: defenses of the models, by the models, for the models. From a global perspective as existing studies (Nguyen et al. 2022; Ozdayi, Kantarcioglu, and Gel 2021; Blanchard et al. 2017; Li et al. 2020a), benign and infected model updates may appear mixed. If we examine model updates in the perspective of individual model update, the nearest ones may have the same purpose as it since both benign and infected updates wish to exclude each other from aggregation. Thus, if we make each model vote for the closest model updates, the benign model updates may get more votes when the benign clients account for the majority.
@@
    Conducting extensive experiments on 5 real-world datasets to demonstrate the superior attack-resistance of Snowball over state-of-the-art (SOTA) defenses when the data are complicatedly non-IID, PDR is not very high and the ratio of attackers to all clients is relatively high. Also, Snowball brings slight impact on the global model accuracy.
@@ 2 Related Work
Thus, there is a strong demand for an approach that can effectively defend backdoor attacks when benign and infected models are scattered without clear boundaries.
@@ 3 Background
where $f(\mathbf{w},\mathcal{D}_{i})\!\!\coloneqq\!\!\frac{1}{|\mathcal{D}_{i}|}\sum_{\xi\in\mathcal{D}_{i},\xi\sim\mathcal{Z}_{i}}\ell(\mathbf{w},\xi)$ is the average loss $\ell$ on data sample $\xi$ of client $i$, where $\xi$ follows distribution $\mathcal{Z}_{i}$, and $\lambda_{i}$ is the weight of client $i$. In each round $t$ of the total $T$ rounds, $K$ ($K\!\leq\!N$) clients are randomly selected as participants. Participant $i$ trains $\mathbf{w}$ to minimize $f$ for $E$ epochs and submit its model update $\Delta\mathbf{w}_{i,t}$ to the server for aggregation. A certain proportion of the participants in each round conduct backdoor attacks, referred as attackers.
@@ 4.1 Overview
To address these challenges, Snowball goes through two election procedures sequentially before aggregation in each round, i.e., bottom-up election and top-down election, as shown in Figure 2. Bottom-up election is designed with insipration of (Shayan et al. 2021; Qin et al. 2023b) which shifts the view from a global perspective to an individual model perspective. It takes the $K$ collected model updates $\mathcal{W}_{t}=\left\{\Delta\mathbf{w}_{i,t}\right\}^{i\in\mathcal{C}_{t}}$ from clients $\mathcal{C}_{t}$ participating round $t$ as the input, and locates a few model updates the least likely to be infected (challenge 1). In it, each model update votes for several ones closest to it, and a few model updates with the most votes are designated as Selectees, denoted by $\widetilde{\mathcal{W}}_{t}\subset\mathcal{W}_{t}$. Such a individual perspective helps to separate benign and infected model updates at a finer granularity (challenge 2).
@@ 4.2 Bottom-up Election
Principle 1 is mentioned in many studies on non-IID problem of FL (Zhao et al. 2018; Fallah, Mokhtari, and Ozdaglar 2020) and widely used by clustering-based personalized FL (Ghosh et al. 2020; Sattler, Müller, and Samek 2020).
@@
Assumption 1 indicates the difference between data distributions of benign clients $i$ and $j$ depends on $\epsilon_{i}$ and $\epsilon_{j}$. When data among clients are IID, $\forall\epsilon$ is a zero distribution. Taken as a whole, benign and infected model updates may not be clearly distinguishable due to the diversity of $\epsilon$. But if assumption 1 and 2 hold, those model updates closer to a benign one are more likely to be benign, as shown in Figure 1. Thus, each model update votes for those closest to them. More supports of Principle 1 are in Appendix A.1.
@@
16:  while $|\widetilde{\mathcal{W}}|$ < $M$ do
@@
It is hard to clearly define “closeness”, so each model update runs K-means independently to guide its voting. For $\Delta\mathbf{w}_{i,t}$, we select $\check{K}-1$ model updates from $\mathcal{W}_{t}-\left\{\mathbf{w}_{i,t}\right\}$ with the largest $\|\Delta\mathbf{w}_{i,t}-\Delta\mathbf{w}_{j,t}\|^{2}$ as the initial centroids of K-means together with a zero vector with the same shape as $\Delta\mathbf{w}_{i,t}$. During implementation, $\check{K}$ is predetermined through Gap statistic (Tibshirani, Walther, and Hastie 2001) on model updates collected in the first round (detailed in Appendix C.2). After clustering, $\check{K}$ clusters are obtained, and $\Delta\mathbf{w}_{i,t}$, as well as the model updates that belong to the same cluster as its, are voted, as shown in the upper right part of Figure 2.
@@ 4.3 Top-down Election
Bottom-up election provides several trusted model updates. However, since benign and infected model updates share certain similarity, K-means, an approach relying on linear distance, cannot deeply mine their differences. To ensure infected model updates excluded, $\check{M}$ has to be small. To avoid that too few model updates included in aggregation such that the convergence of FL is negatively impacted, a VAE (An and Cho 2015) is introduced to learn the patterns of benign model updates by utilizing its powerful nonlinear latent feature representation. Although $\widetilde{\mathcal{W}}_{t}$ provides a few benign patterns, it is still hard to train a VAE, because 1) $|\widetilde{\mathcal{W}}_{t}|$ is too small, and 2) samples in $\widetilde{\mathcal{W}}_{t}$ follow different distributions, causing large reconstruction error. Thus, we shift the focus of detection on the differences between model updates rather than model updates themselves.
@@
If $\Delta\mathbf{w}_{i}^{*}$ is always filtered out, difference between $\Delta\mathbf{w}_{i}^{B}$ and $\Delta\mathbf{w}_{j}$ is expected to have smaller $L_{2}$ norm than that between $\Delta\mathbf{w}_{i}^{*}$ and $\Delta\mathbf{w}_{j}$ as the global model converges.
@@
If infected updates are continually filtered out, $\exists t^{C}<T$ such that after round $t^{C}$,
@@
We experimentally demonstrate it through the average distance among different types of model updates with infected ones filtered out in Figure 3, and leave the theoretical support in Appendix A.2 The average distance between $\Delta\mathbf{w}_{i}^{B}$ and $\Delta\mathbf{w}_{j}^{B}$ is much smaller than that between $\Delta\mathbf{w}_{i}^{B}$ and the others after certain rounds.
@@
Figure 4: Latent features of (a) model updates and (b) differences $\mathbf{d}$ between them outputed by the VAE encoder.
@@
Usually, $\omega$ is an integer close to 0, making the term on the left of (3) usually smaller than 0. Thus, even if few infected models are wrongly included, the distributions of differences between benign and other model updates are easier to learn. We experimentally demonstrate this in Figure 4, showing that learning the differences between model updates can really outperform learning the model updates themselves in terms of distinguishing infected ones. Therefore, we train a VAE $\mathbf{v}$ to learn differences among updates in $\widetilde{\mathcal{W}}_{t}$ by minimizing the loss $J$ by with for $E^{VI}$ epochs on $\mathcal{U}=\left\{\mathbf{u}_{i,j},\ldots\right\}$, where
@@
    $\forall\Delta\mathbf{w}_{j,t}\in\mathcal{W}_{t}-\widetilde{\mathcal{W}}_{t}$, calculate its score $s_{j}=\sum_{\Delta\mathbf{w}_{i,t}\in\widetilde{\mathcal{W}}_{t}}\operatorname{recon}(\mathbf{u}_{i,j},\mathbf{v}(u_{i,j}))$, and then add $M^{E}$ model updates with the lowest scores to $\widetilde{\mathcal{W}}_{t}$ (Comparing model updates in $\mathcal{W}_{t}$ and $\mathcal{W}_{t}-\widetilde{\mathcal{W}}_{t}$).
@@ 5 Experiments
The experiments indend to show: 1) compared with SOTA approaches, Snowball can effectively defend backdoor attacks with complex non-IIDness, a not high PDR and a relatively large attacker ratio. 2) Snowball has comparable accuracy to FedAvg. 3) the VAE is insensitive to related hyperparameters.
@@ 5.1 Datasets and Compared Approaches
The experiments are conducted on five real-world datasets, i.e., MNIST (Deng 2012), Fashion MNIST (Xiao, Rasul, and Vollgraf 2017), CIFAR-10 (Krizhevsky, Hinton et al. 2009), Federated Extended MNIST (FEMNIST) (Caldas et al. 2018) and Sentiment140 (Sent140) (Caldas et al. 2018). They include image classification (IC) and sentiment analysis tasks, and provide non-IID data with Label Distribution Skew, i.e., different $p_{i}(Y)$, and Feature Distribution Skew, i.e., different $p_{i}(X|Y)$, where the latter is even more complex (Tan et al. 2022). These datasets are either already divided into training and test sets, or randomly divided in the ratio of 9:1. We partition MNIST, Fashion MNIST and CIFAR-10 in a practical non-IID way as (Li et al. 2021; Qin et al. 2023a), where data are sampled to 200 clients in Dirichlet distribution with $\alpha=0.5$. FEMNIST contains data from real users and 3,597 of them with more data are selected as clients. Sent140 contains 660,120 users which only hold 2.42 samples averagely, and following (Zawad et al. 2021), we randomly merge these users to form 2,000 distinct clients.
@@
Table 1: Performance (%) of approaches with label distribution skew.
@@
Table 2: Performance (%) of approaches with feature skew.
@@
We compare Snowball with 9 peers, encompassing representative approaches from various categories mentioned in Section 2: 1) Ideal: an imagined ideal approach that filters out all infected updates; 2) FedAvg (McMahan et al. 2017): FL without any defenses; 3) Krum (Blanchard et al. 2017): Byzantine-robust aggregation; 4) CRFL (Xie et al. 2021): certifiable defense based on model ensemble; 5) RLR (Ozdayi, Kantarcioglu, and Gel 2021): an approach with robust parameter-wise learning rate. 6) FLDetector (Zhang et al. 2022): tracing the history model updates to score them; 7) DnC (Shejwalkar and Houmansadr 2021): scoring model updates based on subsets of parameters; 8) FLAME (Nguyen et al. 2022): integrating clustering, weight-clipping and noise-addition; 9) FLIP (Zhang et al. 2023): conducting adversarial learning on clients.

### v2

@@ Abstract
Existing approaches defend against backdoor attacks in federated learning (FL) mainly through a) mitigating the impact of infected models, or b) excluding infected models. The former negatively impacts model accuracy, while the latter usually relies on globally clear boundaries between benign and infected model updates. However, in reality, model updates can easily become mixed and scattered throughout due to the diverse distributions of local data. This work focuses on excluding infected models in FL. Unlike previous perspectives from a global view, we propose Snowball, a novel anti-backdoor FL framework through bidirectional elections from an individual perspective inspired by one principle deduced by us and two principles in FL and deep learning. It is characterized by a) bottom-up election, where each candidate model update votes to several peer ones such that a few model updates are elected as selectees for aggregation; and b) top-down election, where selectees progressively enlarge themselves through picking up from the candidates. We compare Snowball with state-of-the-art defenses to backdoor attacks in FL on five real-world datasets, demonstrating its superior resistance to backdoor attacks and slight impact on the accuracy of the global model.
@@ 1 Introduction
Existing defenses to backdoor attacks in FL are mainly based on a) mitigating the impact of infected models (Bagdasaryan et al. 2020; Sun et al. 2019; Xie et al. 2021; Nguyen et al. 2022; Zhang et al. 2023) or b) excluding infected models based on their deviations (Blanchard et al. 2017; Ozdayi, Kantarcioglu, and Gel 2021; Fung, Yoon, and Beschastnikh 2018; Rieger et al. 2022; Li et al. 2020a; Shejwalkar and Houmansadr 2021; Zhang et al. 2022; Shi et al. 2022; Nguyen et al. 2022). The former may negatively impact global model accuracy (Yu et al. 2021). The latter assumes globally clear boundaries between benign and infected model updates (Zeng et al. 2022). However, backdoor attacks typically manipulate a limited subset of parameters, resulting in the similarity between benign and infected model updates. Besides, the nature of Non-Independent and Identically Distributed (non-IID) data in FL increases diversity among model updates.
@@
Actually, benign and infected model updates are easy to be mixed with complicatedly non-IID data (practical non-IID (Hsu, Qi, and Brown 2019; Huang et al. 2021) and feature distribution skew (Tan et al. 2022)), or with not very high poison data ratio (PDR). We experimentally demonstrate it in Figure 1, where benign and infected updates are mixedly scattered. In such cases, anomaly detections based on linear similarity may not perform satisfactorily. Besides, when facing a relatively high malicious client ratio (MCR), infected model updates are easier to be mistreated as benign ones, however, many existing defenses are only evaluated with MCR $\leq$ 10% (Xie et al. 2021; Ozdayi, Kantarcioglu, and Gel 2021; Zeng et al. 2022; Lu et al. 2022). Although model deviations may be better captured by nonlinear neural networks, the patterns of benign models in FL are usually hard to acquire due to unpredictable distributions and trajectory shifts of model updates. Li et al. (2020a) use the test data to generate model weights for training the detection model, but the test data with a similar distribution to all clients may be usually unavailable. Besides, model weights usually follow extremely complex distributions, making them hard to learn.
To better leverage powerful neural networks to detect malicious models, we propose Snowball, an anti-backdoor FL framework taking advantage of linear and non-linear approaches, i.e., without the need for pre-defined benign patterns and the powerful capability to capture model deviations, respectively. It treats each model update as an agent electing model updates for aggregation with an individual perspective, where the motivation comes from that: defenses of the models, by the models, for the models. From a global perspective as existing studies (Nguyen et al. 2022; Ozdayi, Kantarcioglu, and Gel 2021; Blanchard et al. 2017; Li et al. 2020a), benign and infected model updates may appear mixed. If we examine model updates from the perspective of individual model updates, the nearest ones may have the same purpose since both benign and infected updates wish to exclude each other from aggregation. Thus, if we make each model vote for the closest model updates, the benign model updates may get more votes when the benign clients account for the majority.
@@
    Conducting extensive experiments on 5 real-world datasets to demonstrate the superior attack-resistance of Snowball over state-of-the-art (SOTA) defenses when the data are complicatedly non-IID, PDR is not very high and the ratio of attackers to all clients is relatively high. Also, Snowball brings a slight impact on the global model accuracy. Codes are available at https://github.com/zhenqincn/Snowball.
@@ 2 Related Work
Thus, there is a strong demand for an approach that can effectively defend against backdoor attacks when benign and infected models are scattered without clear boundaries.
@@ 3 Background
where $f(\mathbf{w},\mathcal{D}_{i})\!\!\coloneqq\!\!\frac{1}{|\mathcal{D}_{i}|}\sum_{\xi\in\mathcal{D}_{i},\xi\sim\mathcal{Z}_{i}}\ell(\mathbf{w},\xi)$ is the average loss $\ell$ on data sample $\xi$ of client $i$, where $\xi$ follows distribution $\mathcal{Z}_{i}$, and $\lambda_{i}$ is the weight of client $i$. In each round $t$ of the total $T$ rounds, $K$ ($K\!\leq\!N$) clients are randomly selected as participants. Participant $i$ trains $\mathbf{w}$ to minimize $f$ for $E$ epochs and submit its model update $\Delta\mathbf{w}_{i,t}$ to the server for aggregation. A certain proportion of the participants in each round conduct backdoor attacks, referred to as attackers.
@@ 4.1 Overview
To address these challenges, Snowball goes through two election procedures sequentially before aggregation in each round, i.e., bottom-up election and top-down election, as shown in Figure 2. Bottom-up election is designed with the inspiration of (Shayan et al. 2021; Qin et al. 2024) which shifts the view from a global perspective to an individual model perspective. It takes the $K$ collected model updates $\mathcal{W}_{t}=\left\{\Delta\mathbf{w}_{i,t}\right\}^{i\in\mathcal{C}_{t}}$ from clients $\mathcal{C}_{t}$ participating round $t$ as the input, and locates a few model updates the least likely to be infected (Challenge 1). In it, each model update votes for several ones closest to it, and a few model updates with the most votes are designated as Selectees, denoted by $\widetilde{\mathcal{W}}_{t}\subset\mathcal{W}_{t}$. Such an individual perspective helps to separate benign and infected model updates at a finer granularity (Challenge 2).
@@ 4.2 Bottom-up Election
Principle 1 is mentioned in many studies on the non-IIDness of FL (Zhao et al. 2018; Fallah, Mokhtari, and Ozdaglar 2020) and widely used by clustering-based FL (Ghosh et al. 2020; Sattler, Müller, and Samek 2020).
@@
Assumption 1 indicates that the difference between data distributions of benign clients $i$ and $j$ depends on $\epsilon_{i}$ and $\epsilon_{j}$. When data among clients are IID, $\forall\epsilon$ is a zero distribution. Taken as a whole, benign and infected model updates may not be clearly distinguishable due to the diversity of $\epsilon$. But if Assumptions 1 and 2 hold, those model updates closer to a benign one are more likely to be benign, as illustrated in Figure 1. Thus, each model update votes for those closest to them. More supports of Principle 1 are in Appendix A.1.
@@
16:  while $|\widetilde{\mathcal{W}}|$ $M$ do
@@
It is hard to clearly define “closeness”, so each model update runs K-means independently to guide its voting. For $\Delta\mathbf{w}_{i,t}$, we select $\check{K}-1$ model updates from $\mathcal{W}_{t}-\left\{\mathbf{w}_{i,t}\right\}$ with the largest $\|\Delta\mathbf{w}_{i,t}-\Delta\mathbf{w}_{j,t}\|^{2}$ as the initial centroids of K-means together with a zero vector with the same shape as $\Delta\mathbf{w}_{i,t}$. During implementation, $\check{K}$ is predetermined through Gap statistic (Tibshirani, Walther, and Hastie 2001) on model updates collected in the first round, where the details can refer to in Appendix C.2. After clustering, $\check{K}$ clusters are obtained, and $\Delta\mathbf{w}_{i,t}$, as well as the model updates that belong to the same cluster as its, are voted, as shown in the upper right part of Figure 2.
@@ 4.3 Top-down Election
Bottom-up election provides several trusted model updates. However, since benign and infected updates share certain similarities, K-means, an approach relying on linear distance, cannot deeply mine their differences. To ensure infected updates are excluded, $\check{M}$ has to be small. To avoid too few model updates included in aggregation such that the convergence of FL is negatively impacted, a VAE (An and Cho 2015) is introduced to learn the patterns of benign model updates by utilizing its nonlinear latent feature representation. Although $\widetilde{\mathcal{W}}_{t}$ provides a few benign patterns, it is still hard to train a VAE since 1) $|\widetilde{\mathcal{W}}_{t}|$ is too small, and 2) samples in $\widetilde{\mathcal{W}}_{t}$ follow different distributions, causing large reconstruction error. Thus, we focus on the differences between model updates rather than model updates themselves.
@@
If $\Delta\mathbf{w}_{i}^{*}$ is always filtered out in each round, the difference between $\Delta\mathbf{w}_{i}^{B}$ and $\Delta\mathbf{w}_{j}$ is expected to have a smaller $L_{2}$ norm than that between $\Delta\mathbf{w}_{i}^{*}$ and $\Delta\mathbf{w}_{j}$ as the global model converges.
@@
If infected updates are continually filtered out, $\exists t^{C}<T$ such that after round $t^{C}$, we have
@@
We experimentally demonstrate it through the average distance among different types of model updates with infected ones filtered out in Figure 3. The average distance between $\Delta\mathbf{w}_{i}^{B}$ and $\Delta\mathbf{w}_{j}^{B}$ is much smaller than that between $\Delta\mathbf{w}_{i}^{B}$ and the others after certain rounds. Limited by space, the theoretical support is left in Appendix A.2.
@@
Figure 4: Latent features of (a) model updates and (b) differences $\mathbf{d}$ between them outputted by the VAE encoder.
@@
Usually, $\omega$ is an integer close to 0, making the term on the left of (3) smaller than 0. Thus, even if a few infected updates are wrongly included, the distributions of differences between benign and other updates are easier to learn. Figure 4 experimentally demonstrates that learning the differences between updates outperforms learning the model updates themselves on distinguishing infected ones. Therefore, we train a VAE $\mathbf{v}$ to learn differences among updates in $\widetilde{\mathcal{W}}_{t}$ by minimizing the loss $J$ by with for $E^{VI}$ epochs on $\mathcal{U}=\left\{\mathbf{u}_{i,j},\ldots\right\}$, where
@@
    $\forall\Delta\mathbf{w}_{j,t}\in\mathcal{W}_{t}-\widetilde{\mathcal{W}}_{t}$, calculate its score $s_{j}=\sum_{\Delta\mathbf{w}_{i,t}\in\widetilde{\mathcal{W}}_{t}}\operatorname{recon}(\mathbf{u}_{i,j},\mathbf{v}(u_{i,j}))$, then add $M^{E}$ model updates with the lowest scores to $\widetilde{\mathcal{W}}_{t}$
@@ 5 Experiments
The experiments aim to show: 1) Snowball effectively defends against backdoor attacks with complex non-IIDness, a not high PDR and a relatively large MCR compared to SOTA defenses. 2) Snowball has comparable accuracy to FedAvg. 3) VAE in Snowball is insensitive to hyperparameters.
@@ 5.1 Datasets and Compared Approaches
The experiments are conducted on five real-world datasets, i.e., MNIST (Deng 2012), Fashion MNIST (Xiao, Rasul, and Vollgraf 2017), CIFAR-10 (Krizhevsky, Hinton et al. 2009), Federated Extended MNIST (FEMNIST) (Caldas et al. 2018) and Sentiment140 (Sent140) (Caldas et al. 2018). They include image classification (IC) and sentiment analysis tasks and provide non-IID data with Label Distribution Skew, i.e., different $p_{i}(Y)$, and Feature Distribution Skew, i.e., different $p_{i}(X|Y)$, where the latter is even more complex (Tan et al. 2022). These datasets are either already divided into training and test sets, or randomly divided in the ratio of 9:1. We partition MNIST, Fashion MNIST and CIFAR-10 in a practical non-IID way as (Li et al. 2021; Qin et al. 2023), where data are sampled to 200 clients in Dirichlet distribution with $\alpha=0.5$. FEMNIST contains data from real users and 3,597 of them with more data are selected as clients. Sent140 contains 660,120 users which only hold 2.42 samples averagely, and following (Zawad et al. 2021), we randomly merge these users to form 2,000 distinct clients.
@@
Table 1: Performance (%) of approaches with label distribution skew.
@@
Table 2: Performance (%) of approaches with feature skew.
@@
We compare Snowball with 9 peers, encompassing representative approaches from various categories mentioned in Related Works: 1) Ideal: an imagined ideal approach that filters out all infected updates; 2) FedAvg (McMahan et al. 2017): FL without any defenses; 3) Krum (Blanchard et al. 2017): Byzantine-robust aggregation; 4) CRFL (Xie et al. 2021): certifiable defense based on model ensemble; 5) RLR (Ozdayi, Kantarcioglu, and Gel 2021): an approach with robust parameter-wise learning rate. 6) FLDetector (Zhang et al. 2022): tracing the history model updates to score them; 7) DnC (Shejwalkar and Houmansadr 2021): scoring model updates based on subsets of parameters; 8) FLAME (Nguyen et al. 2022): integrating clustering, weight-clipping and noise-addition; 9) FLIP (Zhang et al. 2023): conducting adversarial learning on clients.